OpenStudy (anonymous):

To which graph does the point (2, -3) belong? HElp please i will give medal to first that answers

OpenStudy (jamesj):

What graphs?

OpenStudy (anonymous):

that is the whole question do you want the multipul choice

OpenStudy (anonymous):

a. y > -x - 2 b. y > -x - 1 c. y > -2x + 3 d. y > -3x + 5

OpenStudy (anonymous):

JamesJ please help

OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, -2) and is parallel to the line y = -4x + 1?

OpenStudy (anonymous):

that is another question

OpenStudy (jamesj):

So to find out of (2,-3) is in any of these regions, test each one in turn: a. y > -x - 2 b. y > -x - 1 c. y > -2x + 3 d. y > -3x + 5 For example, a: If y > -x - 2 is satisfied with (x,y) = (2,-3) then -3 > -2 - 2 = -4 Now that is true as indeed -3 > -4 Test the other three to convince yourself there is only one correct answer here or more than one.

OpenStudy (anonymous):

ya but i don't what i am looking for

OpenStudy (anonymous):

apoorvk help me please, i really need help with this and a few others

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

@JamesJ

OpenStudy (anonymous):

james j can you help with the 1st and 2nd

OpenStudy (anonymous):

@JamesJ

OpenStudy (anonymous):

@JamesJ is the answer c